OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + y = 3 2x + 2y = 10 a. (1, 2) b. (2, 3) c. No Solutions d. Infinitely Many Solutions

OpenStudy (anonymous):

x + y = 3 -y -y x = -y + 3 2(-y + 3) + 2y = 10 -2y + 6 + 2y = 10 -6 -6 y = 4 x + 4 = 3 x = - 1

OpenStudy (anonymous):

i believe there is "no solution"

OpenStudy (anonymous):

you would be correct.

OpenStudy (anonymous):

Thank you!